# Refining Sentence Embedding Model through Ranking Sentences Generation with Large Language Models

### v1

@@ Abstract
Liyang He1,2, Chenglong Liu1,2, Rui Li1,2, Zhenya Huang1,2, Shulan Ruan1,2 Jun Zhou3, Enhong Chen1,2thanks: Corresponding author 1School of Computer Science and Technology, University of Science and Technology of China 2State Key Laboratory of Cognitive Intelligence; 3Ant Group {heliyang,sepcnt,ruili2000,slruan}@mail.ustc.edu.cn, {huangzhy, cheneh}@ustc.edu.cn, jun.zhoujun@antgroup.com
@@ 1 Introduction
Sentence embedding is a fundamental task in natural language processing. It provides effective semantic representations for various downstream applications, such as semantic search He et al. (2023), text classification Wang et al. (2022a), and question-answering systems Nguyen et al. (2022). In recent years, significant progress has been made in the study of sentence embeddings, with methods based on contrastive learning standing out in particular. These approaches learn embeddings of sentences by bringing semantically similar sentences closer together and pushing dissimilar ones further apart. Current mainstream research relies on high-quality annotated data, especially natural language inference (NLI) datasets Bowman et al. (2015); Williams et al. (2018). For instance, supervised contrastive learning methods based on NLI have demonstrated a remarkable ability to surpass the unsupervised approaches Limkonchotiwat et al. (2022); Jiang et al. (2022). However, such annotated datasets are often unavailable in most real-world scenarios, and the manual construction of these datasets incurs extremely high costs.
@@
To reduce reliance on manually annotated data, recent studies have begun to explore leveraging the powerful generative capabilities of large language models (LLMs) to construct high-quality sentence pairs automatically. For instance, SynCSE Zhang et al. (2023) employs LLMs to generate semantically similar sentence pairs, enhancing the effectiveness of contrastive learning. MultiCSR Wang et al. (2024) further evaluates the quality of LLM-generated outputs, filtering out erroneous results. GCSE Lai et al. (2024) utilizes knowledge graphs to extract entities and quantities, enabling LLMs to generate more diverse and knowledge-enriched samples. These approaches significantly diminish the dependence on manual annotation.
@@ 2 Background
Data Generation with LLM. Unsupervised approaches often lag behind their supervised counterparts, which leverage labeled datasets such as natural language inference (NLI) corpora Bowman et al. (2015); Williams et al. (2018). The NLI dataset provides each $x$ with a positive sample $x^{+}$ and a hard negative sample $x^{-}$ to construct the triplet $(x,x^{+},x^{-})$ for the supervised contrastive loss. However, such annotated data are typically unavailable in the majority of scenarios. Thus, researchers began exploring the potential of LLMs for the triplet $(x,x^{+},x^{-})$ generation for each $x\in\mathcal{X}$. A representative work is SynCSE Zhang et al. (2023), which leverages ChatGPT OpenAI (2022) in a few-shot setting to generate positive samples and hard negative samples. MultiCSR Wang et al. (2024) and GCSE Lai et al. (2024) further refined the process of utilizing LLMs for data generation. These works fundamentally revolve around generating triplets.
@@ 3.2 Model Post-training
After obtaining the ranking sentences $l$ for each $x\in\mathcal{X}$, we aim to post-training the existing sentence embedding model to enhance its ability to distinguish fine-grained semantic differences. The ranking sentence $l$ provides order information among sentences. However, the semantic gaps between these sentences are not evenly spaced. Thus, we propose a post-training method that considers both ranking and semantic information among the ranking sentences.
@@
Finally, we post-training the sentence embedding model using the ListMLE Xia et al. (2008) loss. Suppose the representation of $x^{(j)}$ obtained through the sentence embedding model is $\boldsymbol{e}^{(j)}$. Similarly, we can get a similarity relationships list $s^{j}=[s_{j,1},s_{j,2},\dots,s_{j,n}]$. The objective of ListMLE for ranking sentence $l$ is defined as:
@@ 4.1 Dataset
Similar to SynCSE Zhang et al. (2023) and MultiCSR Wang et al. (2024), we utilize the premises of the NLI dataset Bowman et al. (2015); Williams et al. (2018) as the initial unlabeled dataset, denoted as $X_{1}$. Unlike SynCSE and MultiCSR, which employ the full dataset, we sample only a subset for a generation. Specifically, to enhance data diversity, we first apply k-means clustering to $X_{1}$. We set the number of cluster centers to 1,000 and then performed random sampling, selecting 20 samples per cluster, resulting in the dataset $X_{2}$. Next, we generate ranking sentences for each sentence in $X_{2}$ using our method, where the generation step is set to 32, and the hyperparameter $\gamma$ is set to 1.5. In contrast to SynCSE, which relies on ChatGPT with approximately 175B parameters for generation, we utilize the LLaMA3-8B-Instruct. This process ultimately produces the dataset $X_{3}$, consisting of 16,063 sentence ranking lists and a total of 530,079 sentences. Appendix A presents the detailed generation process of our method.
@@ 4.2 Experiment Setup
Baselines. We chose the following strong baselines, including SimCSE Gao et al. (2021), DiffCSE Chuang et al. (2022), PromptBERT Jiang et al. (2022), PCL Wu et al. (2022a), DebCSE Miao et al. (2023), InfoCSE Wu et al. (2022b), RankCSE Liu et al. (2023), SynCSE Zhang et al. (2023), and MultiCSR Wang et al. (2024). Our model is built upon existing sentence embedding models as a post-training approach. In the following experiments, we primarily selected two SOTA models, SynCSE and MultiSCR, as our base models to evaluate whether integrating our ranked sentence data and post-training method can enhance performance [^3] By the time this work was completed, GCSE Lai et al. (2024) was one of the most recent approaches utilizing synthetic data for sentence embedding training. However, as it has not yet publicly released its code and dataset, we did not consider it as a base model.. We designate the post-trained SynCSE and MultiSCR as SynCSE-r and MultiSCR-r, respectively. The details of our training process are provided in Appendix C.
@@ 5 Related Work
To address these limitations, researchers began exploring sentence generation for unlabeled data Chen et al. (2022); Ye et al. (2022) using models like T5 Chung et al. (2024). With the advent of large language models (LLMs), both data annotation and generation have seen significant improvements Gilardi et al. (2023); Alizadeh et al. (2023, 2025). SynCSE Zhang et al. (2023) leverages LLMs to generate semantically similar sentence pairs, enhancing the effectiveness of contrastive learning. MultiCSR Wang et al. (2024) and GCSE Lai et al. (2024) further refine the utilization of LLMs for data generation. This line of research builds upon the training paradigm of supervised SimCSE Gao et al. (2021), where a triplet is generated for contrastive learning. In contrast to these works, our approach shifts the generation objective towards ranking sentence generation, introducing a novel refinement strategy for contrastive learning models.
@@ 7 Limitations
Although our work has achieved a new SOTA performance for existing sentence embedding models, several promising directions still need to be explored, which we leave for future research. In the realm of data synthesis, this paper primarily concentrates on the process of data generation. However, the selection and refinement of synthesized data are equally crucial. For instance, the analysis in An et al. (2024) highlights the issue that synthesized sentences tend to be longer than the original ones. The selection mechanism for synthesized ranking sentence datasets has yet to be explored. Besides, during the post-training process, our primary approach is to integrate ranking information with semantic information. This process involves a hyperparameter $\omega$, whose magnitude influences the model’s performance. Exploring an adaptive method to eliminate the dependence on $\omega$ is also a worthwhile consideration.
@@ Appendix C Model Training Details
All our experimental code is implemented using Python and the PyTorch library. The experiments were conducted on a Linux server equipped with eight NVIDIA GeForce RTX 4090 GPUs. We utilized the official implementations for SynCSE Zhang et al. (2023) and MultiCSR Wang et al. (2024). Specifically, SynCSE provides both model training code and a synthesized dataset. We used their code and dataset to train SynCSE models, including BERT-base, BERT-large, RoBERTa-base, and RoBERTa-large. MultiCSR offers both training and data synthesis code, which we employ to generate data before proceeding with MultiCSR model training. Additionally, in Section 4.6, we reproduce several models, including SimCSE Gao et al. (2021), InfoCSE Wu et al. (2022b), PCL Wu et al. (2022a), and RankCSE Liu et al. (2023). We downloaded their checkpoints from the official HuggingFace repositories and applied our post-training method. During post-training, each model receives a ranking sentence as input per training step. We use the Adam Kingma and Ba (2014) optimizer and set the learning rate to $3\times 10^{-6}$. For SynCSE, the hyperparameter $\omega$ is set to 0.7, while for the other models, $\omega$ is set to 0.5.
@@ Appendix F Transfer Task
Table 7: Comparison of different sentence embedding models accuracy on transfer tasks. The value highlighted in bold is the best value, and the value underlined is the second-best value. $\spadesuit$”: results from Liu et al. (2023). $\clubsuit$”: results from Wang et al. (2024). “*”: we reproduce the results with the officially released corpus from Zhang et al. (2023) and Wang et al. (2024).

### v2

@@ Abstract
Liyang He1, Chenglong Liu1, Rui Li1, Zhenya Huang1,2, Shulan Ruan3 Jun Zhou4, Enhong Chen1thanks: Corresponding author 1State Key Laboratory of Cognitive Intelligence, University of Science and Technology of China; 2 Institute of Artificial Intelligence, Hefei Comprehensive National Science Center; 3Shenzhen International Graduate School, Tsinghua University; 4Zhejiang University {heliyang,sepcnt,ruili2000}@mail.ustc.edu.cn,slruan@sz.tsinghua.edu.cn {huangzhy, cheneh}@ustc.edu.cn, junzhougucas@gmail.com
@@ 1 Introduction
Sentence embedding is a fundamental task in natural language processing. It provides effective semantic representations for various downstream applications, such as semantic search He et al. (2023); Li et al. (2024a), text classification Wang et al. (2022a), question-answering systems Nguyen et al. (2022), and other specialized domain systems Liu et al. (2019); Xue et al. (2024). In recent years, significant progress has been made in the study of sentence embeddings, with methods based on contrastive learning standing out in particular. These approaches learn embeddings of sentences by bringing semantically similar sentences closer together and pushing dissimilar ones further apart. Current mainstream research relies on high-quality annotated data, especially natural language inference (NLI) datasets Bowman et al. (2015); Williams et al. (2018). For instance, supervised contrastive learning methods based on NLI have demonstrated a remarkable ability to surpass the unsupervised approaches Limkonchotiwat et al. (2022); Jiang et al. (2022). However, such annotated datasets are often unavailable in most real-world scenarios, and the manual construction of these datasets incurs extremely high costs.
@@
To reduce reliance on manually annotated data, recent studies have begun to explore leveraging the powerful generative capabilities of large language models (LLMs) to construct high-quality sentence pairs automatically. For instance, SynCSE Zhang et al. (2023) employs LLMs to generate semantically similar sentence pairs, enhancing the effectiveness of contrastive learning. MultiCSR Wang et al. (2024a) further evaluates the quality of LLM-generated outputs, filtering out erroneous results. GCSE Lai et al. (2024) utilizes knowledge graphs to extract entities and quantities, enabling LLMs to generate more diverse and knowledge-enriched samples. These approaches significantly diminish the dependence on manual annotation.
@@ 2 Background
Data Generation with LLM. Unsupervised approaches often lag behind their supervised counterparts, which leverage labeled datasets such as natural language inference (NLI) corpora Bowman et al. (2015); Williams et al. (2018). The NLI dataset provides each $x$ with a positive sample $x^{+}$ and a hard negative sample $x^{-}$ to construct the triplet $(x,x^{+},x^{-})$ for the supervised contrastive loss. However, such annotated data are typically unavailable in most scenarios. Recently, the use of large language models (LLMs) for data construction and annotation has become a trend Li et al. (2024b); Liu et al. (2024). Thus, researchers began exploring the potential of LLMs for the triplet $(x,x^{+},x^{-})$ generation for each $x\in\mathcal{X}$. A representative work is SynCSE Zhang et al. (2023), which leverages ChatGPT OpenAI (2022) in a few-shot setting to generate positive samples and hard negative samples. MultiCSR Wang et al. (2024a) and GCSE Lai et al. (2024) further refined the process of utilizing LLMs for data generation. These works fundamentally revolve around generating triplets.
@@ 3.2 Model Post-training
After obtaining the ranking sentences $l$ for each $x\in\mathcal{X}$, we aim to post-train the existing sentence embedding model to enhance its ability to distinguish fine-grained semantic differences. The ranking sentence $l$ provides order information among sentences. However, the semantic gaps between these sentences are not evenly spaced. Thus, we propose a post-training method that considers both ranking and semantic information among the ranking sentences.
@@
Finally, we post-train the sentence embedding model using the ListMLE Xia et al. (2008) loss. Suppose the representation of $x^{(j)}$ obtained through the sentence embedding model is $\boldsymbol{e}^{(j)}$. Similarly, we can get a similarity relationships list $s^{j}=[s_{j,1},s_{j,2},\dots,s_{j,n}]$. The objective of ListMLE for ranking sentence $l$ is defined as:
@@ 4.1 Dataset
Similar to SynCSE Zhang et al. (2023) and MultiCSR Wang et al. (2024a), we utilize the premises of the NLI dataset Bowman et al. (2015); Williams et al. (2018) as the initial unlabeled dataset, denoted as $X_{1}$. Unlike SynCSE and MultiCSR, which employ the full dataset, we sample only a subset for a generation. Specifically, to enhance data diversity, we first apply k-means clustering to $X_{1}$. We set the number of cluster centers to 1,000 and then performed random sampling, selecting 20 samples per cluster, resulting in the dataset $X_{2}$. Next, we generate ranking sentences for each sentence in $X_{2}$ using our method, where the generation step is set to 32, and the hyperparameter $\gamma$ is set to 1.5. In contrast to SynCSE, which relies on ChatGPT with approximately 175B parameters for generation, we utilize the LLaMA3-8B-Instruct. This process ultimately produces the dataset $X_{3}$, consisting of 16,063 sentence ranking lists and a total of 530,079 sentences. Appendix A presents the detailed generation process of our method.
@@ 4.2 Experiment Setup
Baselines. We chose the following strong baselines, including SimCSE Gao et al. (2021), DiffCSE Chuang et al. (2022), PromptBERT Jiang et al. (2022), PCL Wu et al. (2022a), DebCSE Miao et al. (2023), InfoCSE Wu et al. (2022b), RankCSE Liu et al. (2023), SynCSE Zhang et al. (2023), and MultiCSR Wang et al. (2024a). Our model is built upon existing sentence embedding models as a post-training approach. In the following experiments, we primarily selected two SOTA models, SynCSE and MultiCSR, as our base models to evaluate whether integrating our ranked sentence data and post-training method can enhance performance [^3] By the time this work was completed, GCSE Lai et al. (2024) was one of the most recent approaches utilizing synthetic data for sentence embedding training. However, as it has not yet publicly released its code and dataset, we did not consider it as a base model.. We designate the post-trained SynCSE and MultiCSR as SynCSE-r and MultiCSR-r, respectively. The details of our training process are provided in Appendix C.
@@ 4.5 Analysis
Stage 1 Stage 2
Setting 1 80.91 81.63
Setting 2 77.12 80.74
Table 5: Multi-stage post-training analysis. We used two different post-training sequences and conducted experiments on the STS task, recording the average Spearman’s correlation for each stage.
@@ 4.6 Post-training Experiment
Additionally, we conducted further experiments to examine the impact of multi-stage post-training. Specifically, we utilized the SimCSE-BERT-base model and assessed two different post-training sequences: Setting 1: We first post-trained using the MultiCSR dataset and method, followed by our dataset and method. Setting 2: We began with post-training on our dataset and method, followed by the MultiCSR dataset and method. Table 5 shows these results. In Setting 1, we see that our method improves model performance after post-training with the MultiCSR dataset, achieving a new SOTA performance of 81.63. This highlights the flexibility and compatibility of our approach with different post-training strategies. Moreover, comparing the two settings, MultiCSR offers a greater performance boost when used as the first stage of post-training. This supports our claim that our ranking sentence dataset and method are especially effective in enhancing fine-grained semantic understanding in sentence embedding models.
(a) average length
(b) sentence quality
Figure 4: The analysis of the generated ranking sentences involved calculating the average length (in terms of word count) and the average sentence quality for each index. An index of 0 represents the original sentence.
4.7 Ranking Sentences Analysis
We further designed two analytical experiments for the synthetic dataset. First, a previous study An et al. (2024) reports that sentences synthesized using large language models (LLMs) tend to be longer than the original text. Therefore, we investigated whether the generated ranking sentences also exhibit this characteristic. We calculated the average length of sentence tokens for each ranking index of the generated sentences, and the results are presented in Figure 4(a). The findings indicate that our method produces slightly longer sentences than the originals, but this increase stabilizes, preventing excessive redundancy in subsequent sentences.
Second, we analyzed the quality of the generated sentences. Inspired by the Individual Score method Chen et al. (2023), we employed GPT-4o to evaluate each sentence’s quality on a scale from 0 (worst) to 100 (best). We randomly selected 200 ranking sentences along with their original data for this evaluation. Figure 4(b) displays the results. Compared to the original sentences, the generated sentences demonstrate higher quality. These results suggest that the quality of the generated sentences can be reliably assured.
@@ 5 Related Work
To address these limitations, researchers began exploring sentence generation for unlabeled data Chen et al. (2022); Ye et al. (2022) using models like T5 Chung et al. (2024). With the advent of large language models (LLMs), both data annotation and generation have seen significant improvements Gilardi et al. (2023); Alizadeh et al. (2023); Wang et al. (2024b); Alizadeh et al. (2025). SynCSE Zhang et al. (2023) leverages LLMs to generate semantically similar sentence pairs, enhancing the effectiveness of contrastive learning. MultiCSR Wang et al. (2024a) and GCSE Lai et al. (2024) further refine the utilization of LLMs for data generation. This line of research builds upon the training paradigm of supervised SimCSE Gao et al. (2021), where a triplet is generated for contrastive learning. In contrast to these works, our approach shifts the generation objective towards ranking sentence generation, introducing a novel refinement strategy for contrastive learning models.
@@ 7 Limitations
Although our work has achieved a new SOTA performance for existing sentence embedding models, several promising directions still need to be explored. In the realm of data synthesis, this paper primarily concentrates on the process of data generation. However, the selection and refinement of synthesized data are also crucial. Besides, during the post-training process, our primary approach is to integrate ranking and semantic information. This process involves a hyperparameter $\omega$, whose magnitude influences the model’s performance. Exploring an adaptive method to eliminate the dependence on $\omega$ is also a worthwhile consideration.
Acknowledgments
This research was partially supported by the National Natural Science Foundation of China (Grants No.U23A20319, 62477044), the Fundamental Research Funds for the Central Universities (No.WK2150110038), and the National Key Laboratory of Human-Machine Hybrid Augmented Intelligence, Xi’an Jiaotong University (No. HMHAI-202410).
@@ Appendix C Model Training Details
All our experimental code is implemented using Python and the PyTorch library. The experiments were conducted on a Linux server equipped with eight NVIDIA GeForce RTX 4090 GPUs. We utilized the official implementations for SynCSE Zhang et al. (2023) and MultiCSR Wang et al. (2024a). Specifically, SynCSE provides both model training code and a synthesized dataset. We used their code and dataset to train SynCSE models, including BERT-base, BERT-large, RoBERTa-base, and RoBERTa-large. MultiCSR offers both training and data synthesis code, which we employ to generate data before proceeding with MultiCSR model training. Additionally, in Section 4.6, we reproduce several models, including SimCSE Gao et al. (2021), InfoCSE Wu et al. (2022b), PCL Wu et al. (2022a), and RankCSE Liu et al. (2023). We downloaded their checkpoints from the official HuggingFace repositories and applied our post-training method. During post-training, each model receives a ranking sentence as input per training step. We use the Adam Kingma and Ba (2014) optimizer and set the learning rate to $3\times 10^{-6}$. For SynCSE, the hyperparameter $\omega$ is set to 0.7, while for the other models, $\omega$ is set to 0.5.
@@ Appendix F Transfer Task
Table 8: Comparison of different sentence embedding models accuracy on transfer tasks. The value highlighted in bold is the best value, and the value underlined is the second-best value. $\spadesuit$”: results from Liu et al. (2023). $\clubsuit$”: results from Wang et al. (2024a). “*”: we reproduce the results with the officially released corpus from Zhang et al. (2023) and Wang et al. (2024a).